OpenStudy (anonymous):

which annual rate stats gives an increase of 200% in 20 years?

OpenStudy (amistre64):

im sure there are options we can test out.

OpenStudy (amistre64):

you got a picture of them or can type them up in here it be helpful

OpenStudy (anonymous):

no there isn`t thats the question with no help mesthods the question is really annoying i couldnt solve it any way, the lesson is about log if thats help

OpenStudy (amistre64):

well, there are many formulas that can amount to 200% in 20 years so the question is rather vague

OpenStudy (amistre64):

if its refering to continous compunding rate we can determine it; which might be a good guess since you says it loggy

OpenStudy (amistre64):

100% is to double 200% is to duadruple; or x4 right?

OpenStudy (amistre64):

quadruple even ....

OpenStudy (amistre64):

5 + 5(2.00) = 15 ; might be to triple

OpenStudy (amistre64):

ln(3)/20 would be the time it takes to triple in amount

OpenStudy (amistre64):

not the time, the rate;

OpenStudy (amistre64):

\[A=Pe^{rt}\] \[ln\left(A/P=e^{rt}\right)\] \[ln(A/P)=rt\] \[\frac{ln(A/P)}{t}=r\]

OpenStudy (amistre64):

i just cant recall the 200% jargon. im almost sure it means to triple 100(1+2) = 300 ... yeah, thats it lol

OpenStudy (anonymous):

haha ok ty for the help

OpenStudy (amistre64):

yw, and good luck :)